a particle starts from rest with constant acceleration for 20 s. if it travels a distance y1 in the first 10 s and a distance y2 in the next 10 s then :
a. y2=y1
b. y2=2 y1
c. y2=3 y1
d. y1=4 y2
Answer #1

S = ut + 0.5at^2

y1 = 0 + 0.5*a*10^2

= 50a

speed at the end of 10s

v = u + at

v = 0 + 10a

S = ut + 0.5at^2

y2 = 10a*10 + 0.5*a*10^2

= 150a

y2 = 3y1

so option c) y2 = 3y1 is the correct answer